OVHCloud Network Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#20467 — gra-3ab-a9
Scheduled Maintenance Report for Network & Infrastructure
Completed
Certain links on the gra-3ab-a9 router are used at 100%
capacity. The network team is ready to carry out
migration to new 40G and 100G routers, but these
new routers are not well managed by the backend.
Given the urgency, we will begin the migration anyway
because the quality is not in question. Then the code will be
ended for configurations of IP FO etc.

Update(s):

Date: 2016-09-22 16:00:01 UTC
The migration of the last vlans is completed. Traffic should gradually return to each equipment.

Date: 2016-09-22 15:59:29 UTC
Until now the following ranges have been moved:
151.80.33.0/24
151.80.34.0/24
151.80.35.0/24
151.80.36.0/24
149.202.74.0/24
149.202.75.0/24
149.202.76.0/24
149.202.77.0/24
149.202.82.0/24
149.202.83.0/24
149.202.84.0/24
149.202.85.0/24

we will move 4 new VLANs to the new router couple. For now 50G of traffic has been moved.

Date: 2016-09-22 15:56:28 UTC
We are proceeding to the final part of the migration of 4 vlans.

Date: 2016-09-22 15:56:00 UTC
The IPFO have taken longer to converge than the main IP.

Date: 2016-09-22 15:54:48 UTC
We will migrate new 4 vlans on new router couples.

Date: 2016-09-22 15:54:14 UTC
The last VLANs have been migrated.

Date: 2016-09-22 15:53:50 UTC
The 6th VLAN has been migrated. 4 servers do not correctly re-arp and have not migrated.

Date: 2016-09-22 15:51:49 UTC
Migration of 5th VLAN starting...

Date: 2016-09-22 15:51:17 UTC
We have found a solution to be able to support up to 16 vlans on every couple of new routers. We will resume migration of VLANs one by one with no outage expected (just a cut of 2-3 seconds by the time of the re-arp machines).

Date: 2016-09-22 15:47:52 UTC
Migration is not going as expected. The new routers
are not working as planned and the CPU has been quckly
saturated with 4 vlan (1000 servers).
Normally, the router must endure 3x plus minimum without any problems.
We will change the configuration in order to avoid the CPU saturation.
If this situation persists, we will use another router tomorrow.
We will determine if there is a problem with the
hardware or the software (same hardware and 2 manufacturers).

Date: 2016-09-21 17:04:44 UTC
In order to add a hardware management feature, we will patch the routers one by one.

Date: 2016-09-21 15:59:07 UTC
The 3rd VLAN has been migrated.

Date: 2016-09-21 15:58:48 UTC
The first bang is migrated, the IPs are gradually returning

Date: 2016-09-21 15:58:22 UTC
The configuration is ready to migrate the following 7 vlans, that action will be taken in the next few minutes.

Date: 2016-09-21 15:57:52 UTC
A routing session is not properly functioning on one of the routers. We will restart it.

Date: 2016-09-21 15:56:21 UTC
We will move the first networks to N9. There is no outage expected, but the robots will not be on during this operation.
Posted Sep 21, 2016 - 15:55 UTC